OpenStudy (unklerhaukus):

Find the inverse laplace transform of the arctangent of 1/p \[\mathcal L^{-1}\{\arctan\tfrac1p\}\]

OpenStudy (unklerhaukus):

\[ \newcommand \p \newcommand \p \Lap [1] {\operatorname{\mathcal L}\left\{\rule{0pt}{2.2ex}#1\right\} } \p \dd [1] {\,\mathrm d#1 } \p \de [2] {\frac{\mathrm d #1}{\mathrm d#2} } \p \Lin [1] {\operatorname{\mathcal L}^{-1}\left\{#1\right\} } \p \intl [4] {\int\limits_{#1}^{#2}{#3}\dd{#4} } \begin{align*} \Lin{\arctan\tfrac1p} &=\Lin{\intl p\infty{F(p)}p} =\frac{f(t)}t\\ \\ \arctan\tfrac1p &=\intl p\infty{F(p)}p \\ \de{}p\left(\arctan\tfrac1p\right) &=\de{}p\intl p\infty{F(p)}p \\ \frac1{1+1/p^2}\cdot-1/p^2 &=F(\infty)\cdot(\infty)'-F(p)\cdot(p)' \\ \frac{-1/p^2}{1+1/p^2} &=-F(p) \\ \frac{1}{p^2+1} &=F(p) \\ \\ \Lin{\frac{1}{p^2+1}} &=\Lin{F(p)} \\ \sin(t) &=f(t) \\ \\ \Lin{\arctan\tfrac1p} &=\frac{\sin t}t \end{align*} \]